# Accretion Geometry of the New Galactic Black Hole Candidate AT2019wey in the Hard State

### v1

@@ Abstract
We perform broadband spectral and timing studies of the Galactic low-mass black hole candidate AT2019wey during its 2022 outburst, using quasi-simultaneous observations from NICER, Swift, and NuSTAR. The long-term MAXI light curve, along with the hardness–intensity diagram (HID) derived from NICER data, indicates that the source remained in the hard state and did not switch to the soft state. Spectral modeling using two different model combinations reveals that the broadband spectrum is best described by two distinct Comptonizing regions, associated reflection components, and thermal emission from the disk. The harder Comptonizing region dominates () the total flux and is primarily responsible for the observed reflection features from the distant part of the disk. We find that the accretion disk is truncated at a radius of , while the luminosity is of the Eddington limit. Our spectral results also show consistency in the estimated inner disk radius obtained through two independent methods: modeling the disk continuum and the reflection spectrum. The variability studies imply the presence of intrinsic disk variability, likely originating from an instability in the disk. We also detect hard time lags at low frequencies, possibly arising from the inward propagation of mass accretion rate fluctuations from the outer to the inner regions of the accretion disk. Moreover, an observed deviation of the lag-energy spectrum from the log-linear trend at keV is most likely attributed to thermal reverberation, arising from the reprocessing of hard coronal photons in the accretion disk.
@@ 1 Introduction
AT2019wey (also known as SRGA J043520.9552226, SRGE J043523.3552234) was first discovered as an optical transient by the Asteroid Terrestrial-impact Last Alert System (ATLAS) optical survey in December 2019 (Tonry2019), and was later identified as an X-ray source by the extended ROentgen Survey with an Imaging Telescope Array (eROSITA) and the Astronomical Roentgen Telescope - X-ray Concentrator (ART-XC) instruments onboard the Russian-German Spectrum-Roentgen-Gamma (SRG) observatory in March 2020 (Predehl2021; Pavlinsky2021). Initially, the source was classified as a supernova (Mereminskiy2022A&A...661A..32M), and then considered as a potential BL Lac object (2020ATel13576....1L). However, follow-up observations by 2020ATel13932....1Y, which revealed hydrogen emission lines at zero redshift, suggested that AT2019wey is a Galactic low-mass X-ray binary (LMXB) hosting either a black hole or a neutron star as an accretor. Later on Yao_2021_ApJ_920_120 carried out a multiwavelength study of AT2019wey using observations in 2020 during the same outburst and concluded that the companion star is likely a low-mass () object in a short-period orbit (). The authors also found that the source location on the radio and optical luminosity versus X-ray luminosity diagrams aligns more closely with BHXRBs than with neutron star systems, further supporting its classification as a candidate BHXRB. Using optical and X-ray observations in 2020, Mereminskiy2022A&A...661A..32M also proposed that AT2019wey is a low-mass BHXRB. In addition, a comprehensive X-ray analysis by Yao_2021_ApJ_920_121 during the 2020 outburst showed that the spectral and timing properties of AT2019wey are consistent with those of canonical BHXRBs. Further support came from a spectral study by Feng2022SCPMA..6519512F using NuSTAR data, which reinforced the interpretation of the source as a BHXRB. It is noteworthy that the spectral modeling in all these previous X-ray studies of this source was performed using a single Comptonizing region, which could not constrain the electron temperature or the cut-off energy. Moreover, these studies reported a super-solar iron abundance and suggested that AT2019wey is a low-inclination system hosting a maximally rotating black hole.
@@ 2 Observations
The Galactic BHXRB AT2019wey entered its second outburst phase in 2022, which continued thereafter. Figure 1 shows the long-term MAXI light curve in the keV band, illustrating the onset of this outburst in 2022. In this study, we have analyzed two X-ray observations of the source, taken quasi-simultaneously by Swift/X-ray telescope (XRT), NICER, and NuSTAR in 2022. The start dates of these observations are marked on the MAXI light curve in Figure 1, with detailed information provided in Table 1. The source was also continuously monitored by NICER throughout its outburst from 2022 to 2023. We utilized the NICER observations taken during the gray shaded region starting from the onset of the outbursts in Figure 1 to construct the HID. For this, we first derived the count rates in three energy bands of keV, keV, and keV. The hardness ratio (HR) was then calculated by dividing the count rate in the keV band by that in the keV band. Figure 2 shows the HID derived considering these NICER observations. The shape of the HID does not match the canonical shape during a regular outburst, rather the source remained in either the LHS or HIMS throughout the shaded interval without reaching the soft state.
@@ 2.1 NICER
NICER (Gendreau2016SPIE.9905E..1HG) has observed AT2019wey multiple times during its second outburst that started in 2022. Here, we have used one single NICER observation taken quasi-simultaneously with NuSTAR (see Table 1) for detailed spectral and timing studies. For the NICER data analysis, we employed the standard pipeline NICERDAS, available within HEASoft (v.6.33.2), along with the latest calibration database (CALDB version: xti20240216). We generated cleaned event file using the standard task nicerl2. We then used the standard script nicerl3-spect to extract the source and background spectra, where the ‘3C50’ model was used to estimate the background. During the standard analysis, the nicerl3-spect tool automatically incorporates a systematic of approximately to each channel in the keV band. The task nicerl3-spect also generates the corresponding response matrix file (RMF) and ancillary response file (ARF). Lightcurves were generated using the standard task nicerl3-lc.
2.2 NuSTAR
NuSTAR (Harrison2013ApJ...770..103H) observed AT2019wey twice during its ongoing outburst started in 2022. To analyze the NuSTAR Data, we used the standard analysis software NuSTARDAS (v2.1.2) along with the latest calibration database (CALDB version: 20230613) for both the FPMA and FPMB modules. The standard script nupipeline was employed to process the data. We then derived the source and background spectra using the nuproduct task, using circular regions of radius centered at the source and away from the source, respectively. The corresponding RMFs and ARFs were also generated using the same task.
2.3 Swift/XRT
In this study, we utilized a single archival dataset obtained in the Windowed Timing (WT) mode by Swift/XRT (Burrows2000SPIE.4140...64B; Hill2000SPIE.4140...87H). The level1 data were processed to produce a level2 event file using the standard task xrtpipeline (v.0.13.7) and the latest calibration files. We then employed xselect (v.2.5b) to extract the source and background spectra, selecting circular regions with a radius of , centered at the source position and a region away from the source, respectively. The data were found to be free from pile-up, as the source count rate was well below the threshold value for pile-up in the WT mode. Subsequently, the ARF was generated using the standard task xrtmkarf. Source and background light curves were also extracted using the same circular regions, and the source light curve was then corrected for the background using the lcmath task.
@@ 3.1 Spectral Analysis
Initially, for both observations, we used a powerlaw model modified by the Galactic absorption component tbfeo, which also allows both oxygen and iron abundances to vary. The absorption cross-sections were taken from Verner1996ApJ...465..487V, and the elemental abundances were adopted from Wilms2000ApJ...542..914W. Given the higher spectral quality of the NICER data compared to that of the short-exposure Swift/XRT data, we first modeled Obs2 to constrain the oxygen abundance, while keeping the iron abundance fixed at 0.5 for both the datasets (Zdziarski2024ApJ...967L...9Z). The value of the oxygen abundance was then held fixed during the spectral modeling of Obs1. We also added a multi-color disk blackbody component (diskbb; Mitsuda1984PASJ...36..741M) to our model to account for the thermal emission from the accretion disk. For Obs1, we fixed the inner disk temperature at keV due to the poor data quality, and kept the normalization parameter free to vary. However, this model was unable to describe the spectral data and provided poor fit for both the observations, leaving prominent residuals around the iron ( keV) line and reflection hump ( keV). Figure 3 shows the deviations of the NuSTAR/FPMA spectral data for both the observations from the best-fit model tbfeo*(diskbb+powerlaw). We noticed prominent excesses due to iron emission line and reflection hump. We then replaced the powerlaw component with the nthcomp model, which accounts for the thermal Comptonization of the soft seed photons from the accretion disk in the hot corona. The nthcomp model is parameterized by the photon index () and electron temperature (). We tied the seed photon temperature of the nthcomp component with the inner disk temperature () of the diskbb component. The model constant*tbfeo*(diskbb+nthcomp) provided an improved, though still unsatisfactory, fit for both the observations, with dof and for Obs1 and Obs2, respectively. We then incorporated the ireflect (1995MNRAS.273..837M) component and convolved it with nthcomp to account for the reflection from ionized material. Considering the fact that ireflect is a convolution model, we extended the sample energy range from to keV by using the energies command in XSPEC. We allowed the reflection scaling () factor to vary while keeping all other parameters fixed, including the abundances of iron and other elements at their solar values, the inclination angle at (Yao_2021_ApJ_920_121), the disk temperature at K, and the disk ionization parameter at a high value of . Since ireflect does not consider the emission lines, we included a gaussian component to account for the iron line. As mentioned earlier, we allowed the centroid line energy and width of the iron line to vary freely for Obs2, and then fixed the best-fit values in Obs1. This model constant*tbfeo(diskbb+ireflect*nthcomp+gaussian) improves the fit with dof and for Obs1 and Obs2, respectively.
@@
However, the above model could not constrain the electron temperature (), which was pegging to the highest model-defined value of keV. Besides, significant residuals persisted at keV, suggesting that the coronal structure may not be homogeneous and an additional Comptonizing region is required to fit the broadband spectral data adequately. It is worth mentioning here that Yao_2021_ApJ_920_121 studied AT2019wey during its initial discovery outburst in 2019, when the source was in the LHS. The authors employed a single Comptonizing region in their spectral model but were unable to constrain the coronal electron temperature, which they fixed at keV. Apart from this, Feng2022SCPMA..6519512F also performed X-ray spectral analysis using NuSTAR observations by keeping the electron temperature or cut-off energy in their respective model components fixed at keV or keV, respectively. This further suggests that the structure of the corona may be inhomogeneous, and modeling the corona as a single Comptonizing medium results in the electron temperature pegging at the highest model-defined values. Hence, we incorporated another nthcomp component, independent of the first one, into our model. Similar to the first nthcomp, we considered the seed photon temperature of the second Comptonizing region the same as the disk photons. Other parameters, such as photon index, electron temperature, and normalization of this component, were allowed to vary independently from the first Comptonizing component. Addition of this second nthcomp component provided a substantial improvement in the fit statistics and our model constant*tbfeo(diskbb+ireflect*nthcomp+gaussian+ nthcomp) (hereafter Model-1) now offers a satisfactory and statistically acceptable fit to the broadband spectral data with dof and for Obs1 and Obs2, respectively. The best-fit spectral parameters obtained from Model-1 are listed in Table 2, whereas the best-fit spectral models along with data are shown in Figure 4.
@@ 3.2.1 PDSs
We derived rms-normalized and Poisson noise subtracted PDSs in the keV band using Swift/XRT and NICER observations, and cross-power density spectra (CPDSs) in the keV band using NuSTAR observations. For the Swift/XRT and NICER observations, we first derived lightcurves in the keV band with a time bin size of s, corresponding to the Nyquist frequency of Hz. These lightcurves were then utilized to derive the PDSs using the powspec task, available in HEASoft as a part of the XRONOS package. We also derived the CPDSs from the NuSTAR observations in the keV band with the default time bin size embedded in the HENDRICS package (Bachetti2015ApJ...800..109B). We analyzed all the PDSs and CPDSs for the frequency range of Hz within XSPEC, and quoted the uncertainties in the confidence levels. Neither the PDSs nor the CPDSs exhibit any QPO-like narrow and sharp features. Rather, they can be described well with a combination of two broad Lorentzians required for the band-limited noise (BLNs) components, typically observed in the PDSs of BHXRBs in the LHS (Belloni2002ApJ...572..392B; Wilkinson2009MNRAS.397..666W).
@@ 4.1 Spectral State
The long-term MAXI lightcurve indicates that the source was still in the outburst phase. The HID derived using NICER observations in the shaded interval of the MAXI light curve does not exhibit the canonical q-shaped track typically observed during regular outbursts of low-mass BHXRBs. Variations in the count rate and hardness ratio (HR) in the HID suggest that the source likely underwent a transition to the HIMS toward the end of the shaded interval, and then returned to the LHS again, without passing through the SIMS and the HSS. A likely explanation for the source not reaching the soft state and not exhibiting the canonical q-shaped HID is that it remained in an ongoing outburst phase. Similar to this outburst, the previous discovery outburst of the source in 2019 was reported to be a hard-only or failed outburst, during which the source did not transition to the soft state (Yao_2021_ApJ_920_121). Additionally, hard-only or ‘failed’ outbursts have also been observed in several low-mass BHXRBs such as H 1743–322 (Stiele2021ApJ...914...93S; Sahu2024) and GX 339–4 (Alabarta2021MNRAS.507.5507A; Chand2024ApJ...972...20C). The positions of the observations analyzed in this study on the HID clearly indicate that they were obtained during the LHS.

### v3

@@ Abstract
We perform broadband spectral and timing studies of the Galactic low-mass black hole candidate AT2019wey using quasi-simultaneous NICER, Swift, and NuSTAR observations obtained in 2022. The long-term MAXI light curve, along with the hardness–intensity diagram (HID), indicates that the source remained in the hard state and did not switch to the soft state. Spectral modeling using two different model combinations reveals that the broadband spectrum is best described by two distinct Comptonizing regions, associated reflection components, and thermal emission from the disk. The harder Comptonizing region dominates () the total flux and is primarily responsible for the observed reflection features from the distant part of the disk. We find that the accretion disk is truncated at a radius of , while the luminosity is of the Eddington limit, assuming a black hole mass of and distance of 8 kpc. Our spectral results also show consistency in the estimated inner disk radius obtained through two independent methods: modeling the disk continuum and the reflection spectrum. The variability studies imply the presence of intrinsic disk variability, likely originating from an instability in the disk. We also detect hard time lags at low frequencies, possibly arising from the inward propagation of mass accretion rate fluctuations from the outer to the inner regions of the accretion disk. Moreover, an observed deviation of the lag-energy spectrum from the log-linear trend at keV is most likely attributed to thermal reverberation, arising from the reprocessing of hard coronal photons in the accretion disk.
@@ 1 Introduction
AT2019wey (also known as SRGA J043520.9552226, SRGE J043523.3552234) was first discovered as an optical transient in December 2019 by the Zwicky Transient Facility (ZTF; 2019Masci) survey. The event was subsequently observed independently by the Asteroid Terrestrial-impact Last Alert System (ATLAS; 2018PASP..130f4505T) and Global Astrometric Interferometer for Astrophysics (Gaia; 2016gaia) surveys. Later, it was identified as an X-ray source by the extended ROentgen Survey with an Imaging Telescope Array (eROSITA; Predehl2021) and the Astronomical Roentgen Telescope - X-ray Concentrator (ART-XC; Pavlinsky2021) instruments onboard the Spectrum-Roentgen-Gamma (SRG) observatory in March 2020 (2020ATelMereminskiy). Initially, the source was classified as a supernova based on Gaia observations, and was later considered a potential BL Lac object (2020ATel13576....1L). However, follow-up observations by 2020ATel13932....1Y, which revealed hydrogen emission lines at zero redshift, suggested that AT2019wey is a Galactic low-mass X-ray binary (LMXB) hosting either a black hole or a neutron star as an accretor. Later on Yao_2021_ApJ_920_120 carried out a multiwavelength study of AT2019wey using observations in 2020 during the same outburst and concluded that the companion star is likely a low-mass () object in a short-period orbit (). The authors also found that the source location on the radio and optical luminosity versus X-ray luminosity diagrams aligns more closely with BHXRBs than with neutron star systems, further supporting its classification as a candidate BHXRB. Using optical and X-ray observations in 2020, Mereminskiy2022A&A...661A..32M also proposed that AT2019wey is a low-mass BHXRB. In addition, a comprehensive X-ray analysis by Yao_2021_ApJ_920_121 during the 2020 outburst showed that the spectral and timing properties of AT2019wey are consistent with those of canonical BHXRBs. Further support came from a spectral study by Feng2022SCPMA..6519512F using NuSTAR data, which reinforced the interpretation of the source as a BHXRB. It is noteworthy that the spectral modeling in all these previous X-ray studies of this source was performed using a single Comptonizing region, which could not constrain the electron temperature or the cut-off energy. Moreover, these studies reported a super-solar iron abundance and suggested that AT2019wey is a low-inclination system hosting a maximally rotating black hole.
@@ 2 Observations
The Galactic BHXRB AT2019wey went into its first detected outburst in 2019, which appears to have lasted until early 2022. Figure 1 shows the long-term 40-day binned MAXI light curve of AT2019wey in the 2–20 keV band. After a brief quiescent period of about a month, the source resumed its outburst that continued until early 2025, which we refer to here as the “2022 outburst”. In this study, we analyze two X-ray observations of the source during the 2022 outburst, performed quasi-simultaneously by the Swift/X-ray telescope (XRT), NICER, and NuSTAR. The start dates of these observations are marked on the MAXI light curve in Figure 1, with detailed information provided in Table 1. We also utilized MAXI data from the gray shaded region in Figure 1, covering the full outburst, to construct the HID. The hardness ratio (HR) was calculated as the ratio of the keV to keV count rates, and was plotted against the keV count rate. The resulting HID is shown in Figure 2. For comparison, we overplotted the full ‘q’-shaped HID of GX 339–4, derived from MAXI observations during its complete outburst in 2021. The comparison indicates that the 2022 outburst of AT2019wey was a failed, or hard-only, outburst, with the source remaining in the LHS throughout the entire outburst period.
2.1 NuSTAR
NuSTAR (Harrison2013ApJ...770..103H) observed AT2019wey twice during its outburst started in 2022. To analyze the NuSTAR Data, we used the standard analysis software NuSTARDAS (v2.1.2) along with the latest calibration database (CALDB version: 20230613) for both the FPMA and FPMB modules. The standard script nupipeline was employed to process the data. We then derived the source and background spectra using the nuproduct task, using circular regions of radius centered at the source and away from the source, respectively. The corresponding RMFs and ARFs were also generated using the same task.
2.2 Swift/XRT
Swift observed AT2019wey once quasi-simultaneously with NuSTAR. We utilized this dataset obtained in the Windowed Timing (WT) mode by Swift/XRT (Burrows2000SPIE.4140...64B; Hill2000SPIE.4140...87H). The level1 data were processed to produce a level2 event file using the standard task xrtpipeline (v.0.13.7) and the latest calibration files. We then employed xselect (v.2.5b) to extract the source and background spectra, selecting circular regions with a radius of , centered at the source position and a region away from the source, respectively. The data were found to be free from pile-up, as the source count rate was well below the threshold value for pile-up in the WT mode. Subsequently, the ARF was generated using the standard task xrtmkarf. Source and background light curves were also extracted using the same circular regions, and the source light curve was then corrected for the background using the lcmath task.
@@ 2.3 NICER
NICER (Gendreau2016SPIE.9905E..1HG) observed AT2019wey multiple times during its second outburst that started in 2022. Here, we have used one single NICER observation taken quasi-simultaneously with NuSTAR (see Table 1) for detailed spectral and timing studies. For the NICER data analysis, we employed the standard pipeline NICERDAS, available within HEASoft (v.6.33.2), along with the latest calibration database (CALDB version: xti20240216). We generated cleaned event file using the standard task nicerl2. We then used the standard script nicerl3-spect to extract the source and background spectra, where the ‘3C50’ model was used to estimate the background. During the standard analysis, the nicerl3-spect tool automatically incorporates a systematic of approximately to each channel in the keV band. The task nicerl3-spect also generates the corresponding response matrix file (RMF) and ancillary response file (ARF). Lightcurves were generated using the standard task nicerl3-lc.
@@ 3.1 Spectral Analysis
Initially, for both observations, we used a powerlaw model modified by the Galactic absorption component tbfeo, which also allows both oxygen and iron abundances to vary. The absorption cross-sections were taken from Verner1996ApJ...465..487V, and the elemental abundances were adopted from Wilms2000ApJ...542..914W. Given the higher spectral quality of the NICER data compared to that of the short-exposure Swift/XRT data, we first modeled Obs2 to constrain the oxygen abundance in the tbfeo model. Due to difficulty in constraining the iron abundance, we allowed it to vary over the range . These values of the oxygen and iron abundances were then held fixed during the spectral modeling of Obs1. We also added a multi-color disk blackbody component (diskbb; Mitsuda1984PASJ...36..741M) to our model to account for the thermal emission from the accretion disk. For Obs1, we fixed the inner disk temperature at keV due to the poor data quality, and kept the normalization parameter free to vary. However, this model was unable to describe the spectral data and provided poor fit for both the observations, leaving prominent residuals around the iron ( keV) line and reflection hump ( keV). Figure 3 shows the deviations of the NuSTAR/FPMA spectral data for both the observations from the best-fit model tbfeo*(diskbb+powerlaw). We noticed prominent excesses due to iron emission line and reflection hump. We then replaced the powerlaw component with the nthcomp model, which accounts for the thermal Comptonization of the soft seed photons from the accretion disk in the hot corona. The nthcomp model is parameterized by the photon index () and electron temperature (). We tied the seed photon temperature of the nthcomp component with the inner disk temperature () of the diskbb component. The model constant*tbfeo*(diskbb+nthcomp) provided an improved, though still unsatisfactory, fit for both the observations, with dof and for Obs1 and Obs2, respectively. We then incorporated the ireflect (1995MNRAS.273..837M) component and convolved it with nthcomp to account for the reflection from ionized material. Considering the fact that ireflect is a convolution model, we extended the sample energy range from to keV by using the energies command in XSPEC. We allowed the reflection scaling () factor to vary while keeping all other parameters fixed, including the abundances of iron and other elements at their solar values, the inclination angle at (Yao_2021_ApJ_920_121), and the disk temperature at K. The disk ionization parameter was pegging to a very small value, and therefore we fixed it at . Since ireflect does not consider the emission lines, we included a gaussian component to account for the iron line. As mentioned earlier, we allowed the centroid line energy and width of the iron line to vary freely for Obs2, and then fixed the best-fit values in Obs1. This model constant*tbfeo(diskbb+ireflect*nthcomp+gaussian) improves the fit with dof and for Obs1 and Obs2, respectively.
@@
However, the above model could not constrain the electron temperature (), which was pegged to the highest model-defined value of keV. Besides, significant residuals persisted at keV, suggesting that the coronal structure may not be homogeneous and an additional Comptonizing region is required to fit the broadband spectral data adequately. It is worth mentioning here that Yao_2021_ApJ_920_121 studied AT2019wey during its initial discovery outburst in 2019, when the source was in the LHS. The authors employed a single Comptonizing region in their spectral model but were unable to constrain the coronal electron temperature, which they fixed at keV. Apart from this, Feng2022SCPMA..6519512F also performed X-ray spectral analysis using NuSTAR observations by keeping the electron temperature or cut-off energy in their respective model components fixed at keV or keV, respectively. This further suggests that the structure of the corona may be inhomogeneous, and modeling the corona as a single Comptonizing medium results in the electron temperature pegging at the highest model-defined values. Hence, we incorporated another nthcomp component, independent of the first one, into our model. Similar to the first nthcomp, we considered the seed photon temperature of the second Comptonizing region the same as the disk photons. Other parameters, such as photon index, electron temperature, and normalization of this component, were allowed to vary independently from the first Comptonizing component. Addition of this second nthcomp component provided a substantial improvement in the fit statistics and our model constant*tbfeo(diskbb+ireflect*nthcomp+gaussian+ nthcomp) (hereafter Model-1) now offers a satisfactory and statistically acceptable fit to the broadband spectral data with dof and for Obs1 and Obs2, respectively. The best-fit spectral parameters obtained from Model-1 are listed in Table 2, whereas the best-fit spectral models along with data are shown in Figure 4.
@@ 3.2.1 PDSs
We derived rms-normalized and Poisson noise subtracted PDSs in the keV band using Swift/XRT and NICER observations, and cross-power density spectra (CPDSs) in the keV band using NuSTAR observations. For the Swift/XRT and NICER observations, we first derived lightcurves in the keV band with a time bin size of s, corresponding to the Nyquist frequency of Hz. These lightcurves were then utilized to derive the PDSs using the powspec task, available in HEASoft as a part of the XRONOS package. We also derived the CPDSs from the NuSTAR observations in the keV band with the default time bin size embedded in the HENDRICS package (Bachetti2015ApJ...800..109B). We analyzed all the PDSs and CPDSs for the frequency range of Hz within XSPEC, and quoted the uncertainties in the confidence levels. Neither the PDSs nor the CPDSs exhibit any QPO like narrow and sharp features. Rather, they can be described well with a combination of two broad Lorentzians required for the band-limited noise (BLNs) components, typically observed in the PDSs of BHXRBs in the LHS (Belloni2002ApJ...572..392B; Wilkinson2009MNRAS.397..666W).
@@ 4.1 Spectral State
The long-term MAXI light curve shows that the second outburst of AT2019wey, which started in 2022, persisted until early 2025. Comparison of its HID with that of GX 339–4 during its full outburst in 2021 indicates that the 2022 outburst of AT2019wey was a hard-only or failed outburst, with the source remaining in the LHS throughout the entire outburst period without transitioning to the soft state. Similar to this outburst, the previous discovery outburst of the source in 2019 was reported to be a hard-only or failed outburst, during which the source did not transition to the soft state (Yao_2021_ApJ_920_121). Additionally, hard-only or ‘failed’ outbursts have also been observed in several low-mass BHXRBs such as H 1743–322 (Stiele2021ApJ...914...93S; Sahu2024) and GX 339–4 (Alabarta2021MNRAS.507.5507A; Chand2024ApJ...972...20C).